MCQ Questions For Class 10 Maths Areas Related To Circles Question 11. The radii of two circles are 19 cm and 9 cm respectively. The radius of the circle which has circumference equal to the sum of the circumference of two circles is
(a) 35 cm
(b) 10 cm
(c) 21 cm
(d) 28 cm

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer/ Explanation Answer: d Explaination: Reason: Let the radii of two circles be r1 and r2 and the radius of large circle be r. ∴ r1 = 19 cm, r2 = 9 cm Circumference of two circles = C1+ C2 …(where C = circle) = 2πr1 + 2πr2 = 2π × 19 + 2π × 9 = 38π + 18π = 56π ∴ Circumference of large circle = 56π ⇒ 2πr = 56π ⇒ r = 28 ∴ Radius of large circle = 28 cm